# Can LLMs Simulate Human Behavioral Variability? A Case Study in the Phonemic Fluency Task

### v1

@@ 4.1 Distribution of production frequency
Previous studies have observed that word production in verbal fluency tasks typically follows Zipf’s law, a type of power-law relationship where word frequency is inversely proportional to its rank (e.g., Taler et al., 2020). This distribution is characterized by a small number of words that are produced frequently across participants, followed by a “long tail” of words that are produced infrequently, often appearing only once or twice in the entire dataset. In the current study, human participants’ word production closely followed this expected pattern, with an coefficient of and a strong goodness of fit (). The most frequently produced word by human participants was “fun”, mentioned 34 times, followed by “fire” (30 times) and “fan” (28 times).
@@
In contrast, LLMs, both single and ensembled models, demonstrated a notably different pattern. While all models showed good fit to a power-law distribution ( ranging from to ), their higher coefficients (–) reflect a steeper decline in frequency, with top-ranked words dominating the distribution more heavily than in human data. For instance, O3-mini produced “fast” 83 times, and All-Model Mix generated “fish” 90 times–frequencies substantially higher than those observed in human responses. The Top 5 Models ensemble demonstrated the closest fit to the human distribution patterns (), though still with a steeper coefficient (). Table 3 presents the Zipf’s law goodness-of-fit statistics for all models.
Model | ||
Human | 0.89 | 0.92 |
Claude 3.7 Sonnet | 1.19 | 0.88 |
O3-mini | 1.25 | 0.87 |
[…]
Models with 100+ Word Types | 1.22 | 0.88 |
All-Model Mix | 1.28 | 0.90 |
4.2 Linguistic variables and production frequency
Previous research has shown that verbal fluency performance is influenced by multiple linguistic variables. For instance, Taler et al. (2020) reported that semantic neighborhood size and word frequency were significant predictors of word production in semantic fluency tasks, with words from denser semantic neighborhoods and higher frequency being produced more often. Similarly, Cho et al. (2021) found that phonemic fluency production was influenced by lexical-semantic factors such as word frequency, familiarity, word duration, and age of acquisition.
@@ 4.2 Linguistic variables and production frequency
In this study, we examined the relationship between production frequency and core linguistic variables obtained from the English Lexicon Project (Balota et al., 2007), including word length (number of letters), word frequency (log frequency per million words from SUBTLEX corpus), orthographic neighborhood size (number of words that differ by one letter), phonological neighborhood size (number of words that differ by one phoneme), semantic neighborhood size (number of semantically related words), and age of acquisition (estimated age when a word is typically learned). We first conducted simple correlation analyses between these variables and production frequency.
As shown in Table 4, all linguistic variables except for semantic neighborhood size were significantly correlated with frequency of responses in human data. Word frequency showed the strongest positive correlation (), indicating that high-frequency words were more likely to be produced. Word length and age of acquisition were negatively correlated with production frequency ( and , respectively), suggesting a preference for shorter and earlier-acquired words. Both orthographic and phonological neighborhood sizes, often closely correlated and sometimes used interchangeably in psycholinguistic studies, showed positive correlations, with denser neighborhoods associated with higher production frequency.
@@
Among the LLM models, Claude 3.7 Sonnet most closely approximated the human pattern, with highly similar correlation coefficients across variables. O3-mini followed a similar trend but exhibited generally weaker correlations, particularly for orthographic neighborhood size. The ensemble models followed the same general trend but with progressively weaker correlations, especially for the phonological and orthographic neighborhood variables. Despite differences in correlation strength, the ranking of variable importance was consistent across models: word frequency and age of acquisition showed the strongest correlations, while semantic neighborhood size remained non-significant. This aligns with the phonemic nature of the task, which prioritizes phonological over semantic retrieval processes–a contrast to semantic fluency tasks where semantic factors are more predictive (Taler et al., 2020).
@@
[…]
Models with 100+ Word Types | ∗∗∗ | ∗∗∗ | ∗ | ∗ | ∗∗∗ | |
All-Model Mix | ∗∗∗ | ∗∗∗ | ∗∗∗ | |||
@@
Note. Length = word length; WF = log word frequency (SUBTLEX); Ortho_N = orthographic neighborhood size; Phono_N = phonological neighborhood size; Sem_N = semantic neighborhood size; AoA = age of acquisition. ; ; |
@@
To formally establish the factors that influence production frequency, we conducted stepwise regression analyses with all linguistic variables examined in the correlation analyses above. Full statistical details are provided in Appendix B, but a consistent pattern emerged: linguistic variables explained the most variance in production frequency for human responses (adjusted ), followed closely by Claude 3.7 Sonnet (adjusted ), with substantially less variance explained for O3-mini (adjusted ) and the ensemble models (adjusted –).
Claude 3.7 Sonnet most closely matched the human pattern in both variance explained and predictors retained, with word frequency and age of acquisition emerging as significant in both. Orthographic neighborhood size was uniquely important for humans, while word length–absent from the human model–played a stronger role in Claude and O3-mini. The ensemble models diverged more sharply, with age of acquisition emerging as the dominant–or only–predictor, and word frequency playing a minimal or negligible role. These patterns suggest both shared and divergent influences on word production across systems, with greater unpredictability in the outputs of O3-mini and ensemble models compared to humans and Claude.
5 Network Analysis
@@
Building on our earlier findings, we focused the network analysis exclusively on comparing human and Claude 3.7 Sonnet outputs. Claude consistently exhibited the most human-like patterns–response count, variability, production frequency distribution, and linguistic predictors of word choice. In contrast, O3-mini and the ensemble models showed substantially divergent patterns with lower explained variance and different predictor profiles, suggesting fundamentally different underlying mechanisms driving their word production.
5.1 Network construction approach
@@
To better understand the organizational structure of word production in letter F fluency, we constructed correlation-based networks–a common approach in cognitive science for examining the relationship between concepts (Borodkin et al., 2016; Li and Qiu, in press; Siew and Guru, 2023). Given the different number of nodes in the full networks (476 words for humans, 226 for Claude), direct comparison of network metrics may be confounded by size differences (Borodkin et al., 2016). Therefore, we selected 197 words that were produced by both humans and Claude participants to ensure a fairer comparison of network structure.
@@
The network construction process involved three main steps: (1) creating binary matrices indicating which words each participant produced, (2) calculating word-word cosine similarity matrices based on co-occurrence patterns across participants, and (3) applying a thresholding procedure (TMFG; Massara et al., 2017) to identify meaningful connections between words.
5.2 Structural comparison of human and Claude networks
@@
To complement the network analysis, we conducted a representational similarity analysis (RSA; Nili et al., 2014) comparing the original pairwise similarity matrices that underlie network construction. The Spearman correlation between the lower triangular portions of the human and Claude matrices was modest but significant (, ), providing additional evidence that Claude organizes or retrieves words differently than humans. Taken together with the network metric results, this suggests that while Claude can produce human-like words and respond to similar lexical variables, the global structure and retrieval dynamics it exhibits remain meaningfully distinct from human performance.
6 Discussion
This study evaluated whether LLMs can simulate the behavioral variability observed in human performance on a classic cognitive task: phonemic fluency. Across 34 configurations of state-of-the-art LLMs, we systematically examined the impact of prompt specificity, sampling temperature, model type, and ensemble strategies, comparing model-generated outputs to human F fluency data across three levels of analysis. While some configurations, most notably Claude 3.7 Sonnet at high temperature when prompted with full participant information, produced outputs that broadly resembled human responses in terms of average word count and sensitivity to linguistic features, no model fully captured the scope of individual differences or the structural properties characteristic of human lexical retrieval.
At the participant level, most LLM configurations achieved plausible response counts, particularly when given performance information. However, lexical diversity remained consistently lower than in human participants, and this gap persisted across temperature settings and model types. Models like Claude 3.7 Sonnet produced slightly more diverse responses than others, but even the best-performing configurations failed to reach human-level variation in word types. These findings highlight a persistent tendency toward output rigidity: models tend to repeat the same frequent words across different simulated “participants”.
This consistency persisted even when models were prompted with detailed demographic and performance information, and even when responses were sampled at higher temperatures. The role-play strategy used in prior work (Wang et al., 2025) was conceptually similar to our demographic conditioning, but both approaches appear insufficient for inducing human-like variability in model output. Moreover, ensemble outputs across LLMs, as implemented here, did not recover the distributional diversity observed in human responses. This is likely because most model configurations produced overlapping sets of frequent words. In fact, model ensembles amplified high-frequency core vocabulary and suppressed rare or idiosyncratic responses, further narrowing rather than expanding the response space.
At the item level, further differences emerged in the distribution of word production frequencies. While both human and LLM data followed power-law distributions, the steeper decline in LLM data revealed a stronger bias toward high-frequency words, with less diversity in the “long tail” of infrequent responses. When examining the linguistic factors driving word production, we found some similarities between humans and LLMs, with word frequency and age of acquisition emerging as significant predictors across systems. However, important differences were also evident. For human participants, orthographic neighborhood size played a unique and substantial role, with words having more neighbors being produced more frequently. In contrast, this factor was less influential in LLM outputs, which instead showed greater sensitivity to word length. These patterns suggest that humans may rely more heavily on form-based associations during word retrieval, while LLMs appear to prioritize more surface-level features.
Perhaps most revealing were the network analyses, which uncovered fundamental differences in how humans and LLMs organize lexical retrieval. Human responses exhibited stronger local clustering but weaker global integration, in contrast to Claude 3.7 Sonnet, the most human-like model, which optimized for network-wide efficiency. These structural mismatches underscore a core limitation of current LLMs: they can produce plausible words but fail to approximate the associative dynamics of human memory search.
Taken together, our results have important implications for researchers considering LLMs as substitutes for human participants. While LLMs excel at producing grammatical, contextually appropriate outputs, they do not capture the full spectrum of variability seen in human behavioral data. This limitation is particularly relevant for cognitive tasks where individual differences are of theoretical interest. Researchers who use LLMs as proxies for human behavior should be aware that they may significantly underestimate the true variability in their target population. That said, among the models tested, Claude 3.7 Sonnet consistently emerged as the most human-like, and performed far better than other configurations. Future work could focus on extending this model’s capabilities through more targeted persona modeling, potentially incorporating linguistic and cognitive profiles (e.g., selective reading exposure, task history) to better simulate participant-level differences.

### v2

@@ 4.1. Distribution of production frequency
Previous studies have observed that word production in verbal fluency tasks typically follows Zipf’s law, a type of power-law relationship where word frequency is inversely proportional to its rank, expressed as , where denotes the frequency of a word at rank and is the scaling exponent (e.g., Taler et al., 2020). This distribution is characterized by a small number of words that are produced frequently across participants, followed by a “long tail” of words that are produced infrequently, often appearing only once or twice in the entire dataset.
To estimate , we performed linear regression on the log-transformed rank and frequency values; the slope of the fitted line provides the estimate, and indicates goodness of fit. A larger reflects a steeper decline, indicating that top-ranked words dominate more heavily with less diversity in the long tail. Only model configurations that produced at least unique types were included in this analysis, resulting in configurations alongside the human data (Figure 1).
@@
Human participants’ word production closely followed the expected Zipfian pattern, with and . The most frequently produced word was “fun”, mentioned times, followed by “fire” ( times) and “fan” ( times). In contrast, LLMs demonstrated a notably different pattern. While all models showed adequate fit to a power-law distribution ( ranging from to ), their consistently higher coefficients (–) reflect steeper declines in frequency, with top-ranked words dominating the distribution more heavily than in human data. Among the LLMs, Claude 3.7 Sonnet exhibited the closest to the human distribution (, ), though still notably steeper than the human pattern.
4.2. Linguistic variables and production frequency
@@
Previous research has shown that verbal fluency performance is influenced by multiple linguistic variables, including word frequency, familiarity, word duration, and age of acquisition (Cho et al., 2021; Taler et al., 2020). In this study, we examined the relationship between production frequency and core linguistic variables obtained from the English Lexicon Project (Balota et al., 2007), including word length (number of letters), word frequency (log frequency per million words from SUBTLEX corpus), orthographic neighborhood size (number of words that differ by one letter), phonological neighborhood size (number of words that differ by one phoneme), semantic neighborhood size (number of semantically related words), and age of acquisition (estimated age when a word is typically learned). As with the Zipf analysis, only model configurations with at least unique types were included, resulting in configurations alongside human data. We first conducted simple correlation analyses between these variables and production frequency.
@@
As shown in Table 2, all linguistic variables except for semantic neighborhood size were significantly correlated with production frequency in human data. Word frequency showed the strongest positive correlation (), followed by age of acquisition () and orthographic neighborhood size (). Among the LLMs, Claude 3.7 Sonnet most closely approximated the human pattern, with highly similar correlation coefficients across all variables. o3-mini followed a similar trend but with generally weaker correlations, particularly for orthographic neighborhood size. The remaining models showed progressively weaker alignment, with many losing significance for neighborhood variables while retaining word frequency and age of acquisition as reliable predictors. Semantic neighborhood size remained non-significant across all models, consistent with the phonemic nature of the task.
@@
| GPT-5.2 (High Reasoning) | ∗∗∗ | ∗∗∗ | ||||
| Gemini 2.5 Flash | ∗ | ∗∗∗ | ∗ | ∗ | ∗∗∗ | |
| GPT-5.2 (Medium Reasoning) | ∗∗∗ | ∗∗∗ | ||||
| GPT-3.5 Turbo | ∗∗∗ | ∗∗∗ | ||||
| GPT-5.2 (Low Reasoning) | ∗∗∗ | ∗∗∗ | ||||
| Claude 3.7 Sonnet (Thinking) | ∗∗ | ∗∗∗ | ∗∗ | ∗ | ∗∗∗ | |
| GPT-5 | ∗∗∗ | ∗∗∗ | ||||
| GPT-4o | ∗∗ | |||||
| Claude 4.5 Sonnet | ∗∗∗ | ∗∗∗ | ||||
@@
| Note. Length = word length; WF = log word frequency (SUBTLEX); Ortho_N = orthographic neighborhood size; Phono_N = phonological neighborhood size; Sem_N = semantic neighborhood size; AoA = age of acquisition. ; ; | ||||||
@@
To formally establish the factors that influence production frequency, we conducted stepwise regression analyses with all linguistic variables. Full statistical details are provided in Appendix C, but a consistent pattern emerged: linguistic variables explained the most variance in production frequency for human responses (adjusted ), followed closely by Claude 3.7 Sonnet (adjusted ), with substantially less variance explained for o3-mini (adjusted ). Claude 3.7 Sonnet most closely matched the human pattern in both variance explained and predictors retained, with word frequency and age of acquisition emerging as significant in both. Orthographic neighborhood size was uniquely important for humans, while word length played a stronger role in Claude 3.7 Sonnet and o3-mini.
5. Network Analysis
@@
Building on our earlier findings, we focused the network analysis exclusively on comparing human and Claude 3.7 Sonnet outputs. Claude consistently exhibited the most human-like patterns–response count, variability, production frequency distribution, and linguistic predictors of word choice.
5.1. Network construction approach
To better understand the organizational structure of word production in letter F fluency, we constructed correlation-based networks—a common approach in cognitive science for examining the relationships between concepts (Borodkin et al., 2016; Li and Qiu, 2025; Siew and Guru, 2023). This approach is based on the premise that word co-occurrence patterns in fluency tasks reflect underlying associative relationships: the tendency to produce word given that word was also produced indicates a shared associative link between those items. By aggregating these co-occurrence patterns across participants, it is possible to estimate a network structure where nodes represent words and edges capture the strength of association (Borodkin et al., 2016).
@@ 5.1. Network construction approach
Given the different number of nodes in the full networks ( words for humans, for Claude), direct comparison of network metrics may be confounded by size differences (Borodkin et al., 2016). Therefore, we selected words that were produced by both humans and Claude to ensure a fairer comparison.
@@
The network construction involved three steps: (1) creating binary participant-by-word matrices indicating which words each participant produced, (2) computing word-word cosine similarity matrices based on co-occurrence patterns across participants, and (3) applying the Triangulated Maximally Filtered Graph (TMFG; Massara et al., 2017) algorithm to retain only the strongest associations while preserving network connectivity.
5.2. Structural comparison of human and Claude networks
@@
To complement the network analysis, we conducted a representational similarity analysis (RSA; Nili et al., 2014) comparing the original pairwise similarity matrices that underlie network construction. The Spearman correlation between the lower triangular portions of the human and Claude matrices was modest (, ), providing additional evidence that Claude organizes or retrieves words differently than humans. Taken together with the network metric results, this suggests that while Claude can produce human-like words and respond to similar lexical variables, the global structure and retrieval dynamics it exhibits remain meaningfully distinct from human performance.
6. LLM Ensemble Analysis
The preceding analyses demonstrate that no single LLM configuration captures the full scope of human behavioral variability. A natural question is whether combining outputs from multiple models—each potentially drawing on different architectures, training data, and alignment strategies—could better approximate the distributional diversity observed in human responses. While recent ensemble methods in NLP have primarily focused on improving task accuracy through techniques such as majority voting (Li et al., 2024), our goal is the opposite: to explore whether ensembling can simulate the distributional variability characteristic of human behavior.
6.1. Ensemble construction
To construct ensemble outputs, we used a participant-level random sampling procedure. For each of the human participants, we assembled a pool of eligible LLM outputs from all configurations with , excluding any outlier simulations for that participant (see Appendix B). This resulted in to available LLM outputs per participant, depending on how many configurations produced outliers for that individual.
For each ensemble simulation, we independently and randomly selected one LLM output per participant from their available pool, producing simulated participants from a mixture of models. This procedure was repeated times, yielding ensemble samples.
6.2. Ensemble results
Figure 3 shows the distributions of variability metrics across the ensemble simulations, with human values indicated by the red vertical lines. Across all metrics, the ensemble distributions fell far short of human-level variability. The mean number of unique types across simulations was (, range: –), compared to for human participants. Similarly, mean TTR was (, range: –), well below the human TTR of . The same pattern held for idiosyncratic types and ITTTR.
Notably, the ensemble simulations did not substantially exceed the diversity of the best-performing individual model. Claude 3.7 Sonnet alone produced unique types with a TTR of —values that fall within or near the upper tail of the ensemble distribution for types but above the ensemble range for TTR.
The Zipf analysis of ensemble outputs reinforced this conclusion. Across simulations, the mean coefficient was (, range: –), substantially steeper than the human value of and comparable to individual model values (e.g., Claude 3.7 Sonnet: ; o3-mini: ).
6.3. Vocabulary overlap across models
To understand why ensembling failed to increase diversity, we computed pairwise vocabulary overlap coefficients between all models retained for participant-level analyses (see Appendix D for the full pairwise matrix). The overlap coefficient, defined as the size of the intersection divided by the size of the smaller vocabulary, was remarkably high across the board (mean , ).
Within Anthropic models, overlap was especially pronounced (mean ), with some pairs reaching , indicating that switching between model versions barely changes the word pool. OpenAI models showed somewhat more internal diversity (mean ), though this was largely driven by GPT-3.5 Turbo, an older architecture with the lowest overlap with other OpenAI models.
Even across providers, roughly three-quarters of any model’s vocabulary appeared in every other model. This pervasive lexical overlap explains why combining outputs from diverse models yields minimum gains in diversity: the models are drawing from a largely shared word pool, and ensembling amplifies high-frequency core vocabulary rather than expanding into the long tail of rare or idiosyncratic responses.
7. Discussion and Conclusion
This study evaluated whether LLMs can simulate the behavioral variability observed in human phonemic fluency performance. Across models and configurations, no model—individually or in ensemble—captured the full scope of human variability, despite many producing plausible average response counts and showing sensitivity to similar linguistic features.
A key contribution of this study is its comprehensive coverage of models. Our results reveal that LLMs do not perform equally well on this task, and importantly, newer models are not necessarily better. Within Anthropic, the earlier Claude 3.7 Sonnet substantially outperformed more recent models in lexical diversity, and enabling thinking mode consistently reduced rather than increased variability. These findings suggest that researchers conducting LLM-based simulations of human behavior should evaluate multiple models across different providers rather than defaulting to the latest release, as model selection can substantially influence the conclusions drawn.
Beyond differences in variability, our analyses also revealed divergent retrieval mechanisms. At the item level, orthographic neighborhood size played a unique role for human participants, whereas LLMs showed greater sensitivity to word length, suggesting that humans rely more on form-based associations during phonemic retrieval while LLMs prioritize more surface-level features. The network analyses reinforced this divergence: human responses exhibited stronger local clustering but weaker global integration, whereas Claude 3.7 Sonnet optimized for network-wide efficiency. Together, these structural mismatches highlight a core limitation—LLMs can produce plausible words but fail to approximate the associative dynamics of human memory search.
A fundamental concern motivating our ensemble analysis was that using the same model to simulate multiple participants means those simulated individuals share the same foundational knowledge, making it difficult to generate authentically varied outputs (Wang et al., 2025a). Our ensemble approach addressed this by drawing each simulated participant’s response from a randomly selected model. Yet even this strategy failed: the pairwise vocabulary overlap analysis revealed that models across providers draw from a largely shared word pool, with within-family overlap reaching for Anthropic models. This convergence likely reflects the fact that current LLMs are fundamentally trained on similar internet-scale data using transformer-based architectures, leading to shared biases that amplify high-frequency, prototypical responses at the expense of the rare and idiosyncratic words that characterize human individuality. This pattern raises broader concerns about representational bias: if LLMs systematically suppress low-frequency responses, they risk reinforcing dominant linguistic patterns while marginalizing the diversity inherent in real human populations (Guo et al., 2024).
Among all models tested, Claude 3.7 Sonnet consistently emerged as the most human-like, and performed far better than other configurations. Why this particular model outperforms others, including more recent models from the same provider, remains an open question. Future work could focus on extending this model’s capabilities through more targeted persona modeling, potentially incorporating linguistic and cognitive profiles (e.g., selective reading exposure; Qiu et al., 2024) to better simulate participant-level differences. More broadly, rather than substituting LLM-simulated individuals for real human participants, a more productive approach may be to treat LLMs as a baseline system–one that captures central tendencies in linguistic knowledge–against which the uniqueness and flexibility of real human behavior can be measured.